The formula N = 2x^2+6x+3 represents the number of households N, in thousands, in a certain city that have a computer x years after1990. According to the formula, in what year were there 59 thousand households with computers in this city? ( for this problem ignore any negative solutions.)
Answer by jorel1380(3719)

59=2x2+6x+3
2x2+6x-56=0
x2+3x-28=0
(x+7)(x-4)=0
x=-7 or 4
Throwing out the negative answer, we get x=4, so the year was 1990 + 4=1994..